Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by hey_thr67 » Tue Jun 26, 2012 10:53 am
If Ben were to lose the championship, Mike would be the winner with a probability of 1/4, and Rob 1/3. If the probability of Ben being the winner is 1/7, what is the probability that either Mike or Rob will win the championship?

A: 1/12
B: 1/7
C: 1/2
D: 7/12
E: 6/7
Join the discussion
Source: — Problem Solving |

by eagleeye » Tue Jun 26, 2012 2:14 pm
This is a combined probability problem.

Let Probability of Ben, Mike and Rob winning be P(B), P(M), P(R) respectively.

Then we are told that, when Ben doesn't win, the probability of Mike winning is 1/4 and that of Rob is 1/3. Now probability of Ben losing = P(#B) = 1-P(B). (where I have used #B to show the event of Ben not winning. Remember that probability of one winning and the other two also winning is 0.
The event that Ben loses and Mike wins is denoted by (#BM). Now P(#BM) = P(#B)*P(M).

So, we are given that
P(#B)*P(M) = 1/4
P(#B)*P(R) = 1/3

Then we are told that P(B) = 1/7. So P(#B) = 1-1/7 = 6/7.

We need to find probability of either Mike or Rob winning. So we need P(M)+P(R).

In the first two equations, substituting P(#B) = 6/7 and adding them, we get
6/7*(P(M)+P(R)) = 1/4 + 1/3 => P(M)+ P(R) = (1/3+1/4)*6/7 = 7/12*6/7 = 1/2.

Hence C is the right answer.

Let me know if this helps :)
Join the discussion

by Mike@Magoosh » Tue Jun 26, 2012 2:21 pm
hey_thr67 wrote:If Ben were to lose the championship, Mike would be the winner with a probability of 1/4, and Rob 1/3. If the probability of Ben being the winner is 1/7, what is the probability that either Mike or Rob will win the championship?

A: 1/12
B: 1/7
C: 1/2
D: 7/12
E: 6/7
Hi, there. I'm happy to help with this. :-)

This is a tricky question. First of all, P(Ben wins) = 1/7, so P(Ben doesn't win) = 1 - 1/7 = 6/7.

The wording of the first part is very important: If Ben were to lose the championship, Mike would be the winner with a probability of 1/4, and Rob 1/3. This is a conditional probability --- we are told the probability given that a certain condition (Ben losing) is met.

The conditional probability of A, given B, is written P(A|B). The formula for this is

P(A and B) = P(B)*P(A|B)

Here, B = Ben loses, and we want A = Mike or Rob wins.

P(Mike wins|Ben loses) = 1/4
P(Rob wins|Ben loses) = 1/3

Add these two:
P(Mike wins|Ben loses) + P(Rob wins|Ben loses) = P(Mike or Rob wins|Ben loses) = 1/4 + 1/3 = 7/12

P(Mike or Rob wins) = P(Ben loses)*P(Mike or Rob wins|Ben loses) = 6/7 * 7/12 = 6/12 = 1/2

Answer = C

Again, that was a very hard question. This is at the very outer limit of what the GMAT could possible ask. Here's a much more typical kind of GMAT math probability question.

https://gmat.magoosh.com/questions/1036
When you submit your answer to this question, the next page will have the full video solution.

Let me know if you have any more questions.

Mike :-)
Magoosh GMAT Instructor
https://gmat.magoosh.com/
Join the discussion

by GMATGuruNY » Tue Jun 26, 2012 3:21 pm
hey_thr67 wrote:If Ben were to lose the championship, Mike would be the winner with a probability of 1/4, and Rob 1/3. If the probability of Ben being the winner is 1/7, what is the probability that either Mike or Rob will win the championship?

A: 1/12
B: 1/7
C: 1/2
D: 7/12
E: 6/7
When we want one event AND another event to happen -- P(A AND B) -- we MULTIPLY the fractions.
When we want one event OR another event to happen -- P(A OR B) -- we ADD the fractions.

Since P(Ben wins) = 1/7, P(Ben loses) = 6/7.

Case 1: Ben loses AND Mike wins.
Since we want both events to happen, we multiply:
6/7 * 1/4 = 3/14.

Case 2: Ben loses AND Rob wins.
Since we want both events to happen, we multiply:
6/7 * 1/3 = 2/7.

Since either Case 1 OR Case 2 will yield a good outcome, we add the results above:
3/14 + 2/7 = 3/14 + 4/14 = 7/14 = 1/2.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion